OpenStudy (anonymous):

Put into standard form. x^2 + 8x - 10y - 34 = 0 (parabola)

OpenStudy (anonymous):

I have gotten to this. (x+4)^2 = 10y+18. im not sure what o do next

OpenStudy (anonymous):

\[y=0.1x ^{2}+0.8x-3.4\]

OpenStudy (anonymous):

im looking for the form (x-h)^2 = 4p (x-h)

OpenStudy (anonymous):

ooooooh ,yes

OpenStudy (anonymous):

\[(x+4)^{2}=10(y+1.8)\]

OpenStudy (anonymous):

Thats what i got, but what would be the vertex?

OpenStudy (anonymous):

(-4,-1.8) i would assume. If thats rigtht then none of my options are correct :/

OpenStudy (anonymous):

sorry is not correct \[(x+4)^{2}=10(y+4.2)\] is it true?

OpenStudy (anonymous):

the options are (-4,-5) (-4,5) (-4,3.4)

OpenStudy (anonymous):

im sorry i Mixed up .... \[(x+4)^{2}=10(y+5)\] i sure that is right :)

OpenStudy (anonymous):

Ok thank you. Could you please tell me how you got there? just a quick explanation

OpenStudy (anonymous):

@fazeli u a girl ;)

OpenStudy (anonymous):

nice and cute name ;)

OpenStudy (anonymous):

haye itna na tarsao jaldi type karo my fazeli ;)

OpenStudy (anonymous):

thanks you started the question, i only Simplify Phrase.

OpenStudy (anonymous):

thanks anyway..

OpenStudy (whpalmer4):

\[x^2 + 8x - 10y - 34 = 0\]Move all terms not containing \(x\) to the right side to get them out of the way while we complete the square: \[x^2+8x = 10y+34\]Complete the square by adding half the coefficient of the \(x\) term, squared, to both sides: \[x^2+8x+(\frac{8}{2})^2 = 10y+34 + (\frac{8}{2})^2\]\[x^2+8x + 16 = 10y+34 + 16\]Rewrite left hand side as a perfect square:\[(x+4)^2 = 10y+34+16\]\[(x+4)^2 = 10y+50\]Factor right hand side\[(x+4)^2 = 10(y+5)\]\[10=4p\]\[p=\frac{5}{2}\]Standard form is \[(x+4)^2 = 10(y+5),\text{ with }p=\frac{5}{2}\] Standard form gives us \[(x-h)^2 = 4p(y-k)\] \[x+4 = x-h\]\[h=-4\]\[y+5=y-k\]\[k=-5\]Vertex is at \((h,k) = (-4,-5)\) You can check this result by putting equation in vertex form: \[y = a(x-h)^2+k\]\[(x+4)^2 = 10(y+5)\]\[(x+4)^2=10y+50\]\[10y=(x+4)^2-50\]\[y=\frac{1}{10}(x+4)^2-5\]Again, \(h = -4, \,k = -5\) so our original vertex of \((-4,-5)\) was correct.

OpenStudy (whpalmer4):

Here's a plot of the parabola between \(x=-15\) and \(x = 10\) showing the vertex at \((-4,-5)\)

OpenStudy (anonymous):

THANK YOU SO MUCH :D!!! fan

OpenStudy (anonymous):

@whpalmer4 U a girl!!! =D ?

OpenStudy (whpalmer4):

uh, no :-) what gives you that idea?

OpenStudy (anonymous):

HahA :D

OpenStudy (whpalmer4):

I'd make quite the attractive girl — I can cover my eyes with my beard :-)

OpenStudy (anonymous):

WAH WAH :D what a special type of beard!!YOU GO BRO!!!

OpenStudy (whpalmer4):

I'm saving a bundle not buying razor blades :-)

OpenStudy (whpalmer4):

a few hundred years of razor blade savings and I'll be able to buy a used Harley to go with the beard :-)